A metal rod of 10mm square cross section and length 2 m is loaded in tension with an axial tensile load of 4 kN. If the tensile modulus is 72 GPa, what will be the extension of the rod?
1
Expert's answer
2019-08-02T10:23:11-0400

Deformation of bar is given by

"\Delta l=\frac{P\times l}{A\times E}"

Where P=4kN=4x1000N

L=2m=2000mm

A=10x10mm2

E=72GPa=72000MPa=72000N/mm2

Δl=(4000x2000)/(10x10x72000)

=1.11mm